OpenStudy (anonymous):

Simplify 4x/9+(x-10)/9

OpenStudy (anonymous):

\[\frac{ 4x }{ 9 }+\frac{ x-10 }{ 9 }\]

OpenStudy (anonymous):

4x+x-10=9 5x-10=9 5x=19 x=19/5

OpenStudy (anonymous):

could you explain to me what you did, if you dont mind.

OpenStudy (anonymous):

steps by taking lcm multipling both sides by 9 adding 10 to both sides dividing both sides by 5

OpenStudy (anonymous):

\[\frac{ 4x }{ 9 }+\frac{ x-10 }{ 9 }=\frac{ 4x+x-10 }{ 9 }=\frac{ 5x-10 }{ 9 }\]

OpenStudy (anonymous):

@bridgetolivares does that help?

OpenStudy (anonymous):

yes, thanks

OpenStudy (anonymous):

:)